OpenStudy (anonymous):

A toy company manufactures two types of dolls, A and B. Market tests and available resources have indicated that the combined production level should not exceed 1200 dolls per week and the demand for dolls of type B is at most half of that for dolls of type A. Further, the production level of dolls of type A can exceed three times the production of dolls of other type by at most 600 units. If the company makes profit of Rs 12 and Rs 16 per doll respectively on dolls A and B, how many of each should be produced weekly in order to maximize the profit?

OpenStudy (anonymous):

@satellite73 @UnkleRhaukus @karatechopper @Hero

OpenStudy (anonymous):

"indicated that the combined production level should not exceed 1200 dolls per week" \[A+B\leq 1200\]

OpenStudy (anonymous):

\[A\leq 3B+600\]

OpenStudy (anonymous):

sorry i am unable to view it

OpenStudy (anonymous):

\[B\leq \frac{1}{2}A\]

OpenStudy (anonymous):

Those are your restrictions and you need to maximize \[12A+16B=P\]

OpenStudy (anonymous):

can you please tell me how to draw the graph

OpenStudy (anonymous):

@Outkast3r09

hero (hero):

Looks like @Outkast3r09 provided the equations you need

OpenStudy (anonymous):

i know how to do the equations.

hero (hero):

So what do you need help with?

OpenStudy (anonymous):

@Hero its the graph. but i am unable to find its intersection point. help me with that plz

hero (hero):

You want to maximize 12A + 16B = P while adhering to the constraints. So whichever point yields the maximum profit will be your solution.

OpenStudy (anonymous):

i know but before that .i'm unable to find that maximum point :(

hero (hero):

I found it in two seconds

OpenStudy (anonymous):

tell me the process please

hero (hero):

Well, basically, all you do is define your profit equation as f(a,b) = 12a + 16b then try different points until you find the one that yields the maximum profit So try: f(600, 0) f(800, 400) f(1050, 150) One of those will yield the max.

OpenStudy (anonymous):

the points at each corner of the shaded area is the possible max profits

OpenStudy (anonymous):

so the most simple way would to find out at which what points do those lines cross and then plug them into your profit equation

OpenStudy (anonymous):

after that i have to find their intersection point,which i m unable to do

OpenStudy (anonymous):

pull the inequality signs out and you'll get the lines, find the intersects by equaling them to each other

OpenStudy (anonymous):

from the picture you can see what lines intersect

OpenStudy (anonymous):

at c and b

OpenStudy (anonymous):

I mean that if you graphed the lines you could see which functions intersect each other... then make those functions equal each other and you'll get the x coordinate of the point of intersection